# Semantic-Enhanced Time-Series Forecasting via Large Language Models

### v2

@@ Introduction
Current most LLM-based methods mainly focus on token-level alignment (Hu et al. 2025), as shown in Fig. 1 (a) (Pan et al. 2024a; Liu et al. 2025). The joint space serves as an implicit prompt, guiding the LLM in conducting time series forecasting. However, the semantic tokenization eliminates the dependency on auxiliary textual descriptions, leading to noise and computational costs (Jin et al. 2024). To address these limitations, our analysis identifies a critical oversight in existing LLM-based approaches (Jin et al. 2024; Liu et al. 2024b; Chang, Peng, and Chen 2025; Zhou et al. 2023). Regarding the semantic space composed of tokens and time steps, we propose the Temporal Semantic Cross-Correlation (TSCC) Module to obtain enhanced semantics that represent temporal patterns, emphasize anomaly and de-anomaly patterns, thereby enhancing the interpretability of the token embeddings, as shown in Fig. 1 (b).
Although we provide semantic guidance with high interpretability from LLMs, there exists an inherent difference between LLMs’ understanding of linguistic knowledge and time series data, making the modal gap difficult to bridge. In summary, we propose Semantic-Enhanced LLM, a novel framework that enhances LLM for time series forecasting with two key components: the Temporal-Semantic Cross-Correlation (TSCC) module and the Time-Adapter architecture. The TSCC captures cross-correlation of temporal embedding and semantic space with temporal dependencies and channel-wise relationships, which embed into token embedding to adapt LLM to temporal features. The Time-Adapter enhances forecasting capability by overcoming the limitation of LLMs in modeling both short-term and long-term dependencies based on historical time segments. This enables pretrained LLMs to forecast time series effectively. Our main contributions are four-fold:
@@ Methods
Figure 2: (a) An overview of SE-LLM. Red indicates trainable, while blue indicates frozen. (b) The overall framework of TSCC, with detailed processes referring to Fig. 3. (c) For the Transformer-based LLM, the Time-Adapter is embedded within the key and value vectors of the multi-head attention mechanism.
Semantic Space. The semantic space is the semantic embedding obtained by mapping the word embedding of a large language model through a linear layer, which we collectively refer to as the semantic space. Specifically, we map the Word Embedding layer from $l_{1}\in\mathbb{R}^{L_{1}\times C}$ to $l_{2}\in\mathbb{R}^{K\times C}$ (Fang et al. 2025).
The pre-trained word embeddings from LLM are used as general features, and a linear transformation is applied to transfer the features to a new semantic space, allowing the model to match the time series of the TS Embeddings. We input both the semantic space and features into the TSCC to obtain enhanced semantics, as shown in Fig. 2(b). Following this, the LLM embeds Time-Adapter into the Multi-Head Attention to further refine the model’s ability to understand temporal patterns in the time series data, as shown in Fig. 2(c). This component enhances the overall temporal pattern recognition by integrating additional contextual information, enabling better forecasting performance. Finally, the Decoder component decodes the output from the LLM to produce the final prediction, which can be formulated as:
@@
where $\mathbf{O}\in\mathbb{R}^{B\times N\times K}$, $\mathbf{Y}\in\mathbb{R}^{B\times N\times C}$ is the LLM’s output.
@@
The Modified VAE Module. In time series analysis, non-stationary information refers to the time-varying statistical properties of the sequence. Non-stationarity lead to inaccurate model parameter estimation and increased prediction bias. Correctly identifying and handling non-stationarity is a crucial step in time series modeling, directly affecting the accuracy and reliability of predictions.
We employ a Variational Autoencoder (VAE) (Cemgil et al. 2020) to reconstruct the mean and variance in a joint space, comprehensively capturing the statistical characteristics of the sequences. Specifically, we utilize the VAE module to model the underlying stochastic process of noise in temporal data, which refer to Algorithm 1, where $\mathbf{DC}$ represents the latent noise component separated from the joint space, $\mathbf{DA}$ is the denoising component, which contains information such as periodic or trends.
Algorithm 1 Variational Autoencoder (VAE) Module
0: Cross-attention output $\mathbf{C}\in\mathbb{R}^{B\times N\times C}$
0: Decomposed anomaly $\mathbf{DC}$ and de-anomalized component $\mathbf{DA}$
2:Latent Distribution Estimation: $\boldsymbol{\mu},\log\boldsymbol{\sigma}^{2}\leftarrow\mathcal{F}_{\mu}(\mathbf{V}),\mathcal{F}_{\sigma}(\mathbf{V}),$ where $\mathcal{F}_{\mu},\mathcal{F}_{\sigma}$: Linear mapping from $\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$
3:Reparameterization Trick: Sample from latent space, $\mathbf{z}\leftarrow\boldsymbol{\mu}+\boldsymbol{\epsilon}\odot\exp\left(0.5\cdot\log\boldsymbol{\sigma}^{2}\right),\quad\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$
4:Reconstruction: Decode latent variable to recover input, $\hat{\mathbf{C}}\leftarrow\mathcal{F}_{d}(\mathbf{z}),\quad\mathcal{F}_{d}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}\rightarrow\mathbb{R}^{C}$
5:Anomaly Decomposition: $\mathbf{DC}=\mathbf{C}-\hat{\mathbf{C}},\mathbf{DA}=\hat{\mathbf{C}},\mathbf{DC}\in\mathbb{R}^{B\times N\times C},\mathbf{DA}\in\mathbb{R}^{B\times N\times C}$
@@
Subsequently, the cross-correlation matrix $\mathbf{M}\in\mathbb{R}^{B\times K}$ is computed via standardization over both temporal and semantic features, which can be formulated as:
@@
Followed by a top-K (Liu, Zeng, and Ren 2024) filtering mechanism to select the most strongly correlated positions and weighted to the output of VAE to endow both anomaly and de-anomaly semantic regions with a temporal TS Embeddings that exhibits strong correlations. It can be formulated as:
@@
where $\mathbf{\overline{DC}}$ represents enhanced anomaly semantics reconstructed by the VAE module from the joint space, $\mathbf{\overline{DA}}$ represents enhanced de-anomaly semantics, the $\mathbf{\overline{DA}},\mathbf{\overline{DC}}\in\mathbb{R}^{B\times N\times C}$. $\mathcal{P}$ represents the matrix in the semantic space corresponding to the indices of the top-k positions in the cross-correlation matrix.
Channel Dependency Enhancement. The aforementioned enhanced semantics are learned by integrating the TS Embeddings of time series data. To prevent the loss of feature information, we further acquire channel dependencies based on the enhanced semantics modeled along the temporal dimension. Specifically, we concencate the TS Embeddings and the $\overline{\textbf{DA}},\overline{\textbf{DC}}$, that share information across channel dimensions. We employ a Multi-Layer Perceptron (MLP) for attention extraction, perform channel enhancement on the enhanced semantics in the channel dimension, which can be formulated as :
@@
$\mathbf{GA}=\mathrm{Attn_{a}}\odot\mathrm{S^{T}}+(1-\mathrm{Attn_{a}})\odot\mathrm{H}.$ (8)
@@
Models SE-LLM TimeXier++ TimeMOE CSFormer Time-CMA LLM-TS AutoTimes-GPT2 Time-LLM iTransformer DLinear
2025 2025 2025 2025 2025 2024 2024 2024 2023
Datasets Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
96 0.352 0.393 0.375 0.404 0.349 0.379 0.372 0.394 0.366 0.396 0.403 0.420 0.360 0.397 0.398 0.410 0.387 0.418 0.375 0.399
192 0.378 0.411 0.415 0.436 0.395 0.413 0.420 0.425 0.401 0.420 0.440 0.441 0.391 0.419 0.451 0.440 0.416 0.437 0.405 0.416
ETTh1 336 0.391 0.420 0.442 0.451 0.447 0.453 0.453 0.440 0.412 0.431 0.471 0.457 0.408 0.432 0.473 0.451 0.434 0.450 0.439 0.443
720 0.402 0.437 0.482 0.482 0.457 0.462 0.470 0.470 0.406 0.427 0.503 0.487 0.429 0.452 0.469 0.470 0.447 0.473 0.472 0.490
Avg 0.381 0.415 0.429 0.443 0.412 0.426 0.429 0.432 0.396 0.419 0.454 0.451 0.397 0.425 0.448 0.443 0.421 0.445 0.423 0.437
96 0.150 0.200 0.150 0.202 0.198 0.288 0.168 0.217 0.161 0.205 0.166 0.217 0.158 0.208 0.149 0.200 0.174 0.225 0.152 0.237
192 0.194 0.243 0.196 0.244 0.235 0.312 0.213 0.257 0.199 0.249 0.229 0.269 0.207 0.264 0.195 0.243 0.227 0.268 0.220 0.282
Weather 336 0.247 0.285 0.248 0.285 0.293 0.348 0.272 0.298 0.267 0.2900 0.278 0.302 0.262 0.298 0.245 0.282 0.290 0.309 0.265 0.319
720 0.323 0.339 0.316 0.332 0.427 0.428 0.346 0.347 0.343 0.341 0.354 0.351 0.342 0.353 0.318 0.338 0.374 0.360 0.323 0.362
Avg 0.229 0.267 0.228 0.266 0.288 0.344 0.250 0.280 0.243 0.271 0.257 0.285 0.242 0.281 0.227 0.266 0.266 0.291 0.240 0.300
96 0.350 0.243 0.367 0.266 —— —— —— —— —— —— 0.587 0.315 0.369 0.257 0.373 0.280 0.363 0.259 0.410 0.282
192 0.373 0.254 0.379 0.364 —— —— —— —— —— —— 0.612 0.326 0.394 0.268 0.390 0.288 0.383 0.267 0.423 0.287
Traffic 336 0.390 0.263 0.417 0.301 —— —— —— —— —— —— 0.634 0.338 0.413 0.278 0.407 0.299 0.396 0.275 0.436 0.296
720 0.429 0.285 0.453 0.314 —— —— —— —— —— —— 0.640 0.351 0.449 0.299 0.438 0.310 0.435 0.296 0.466 0.315
Avg 0.386 0.261 0.404 0.311 —— —— —— —— —— —— 0.618 0.333 0.406 0.276 0.402 0.294 0.394 0.274 0.434 0.295
96 0.130 0.225 0.132 0.226 0.157 0.211 0.146 0.242 0.151 0.260 0.167 0.271 0.140 0.236 0.137 0.244 0.133 0.229 0.153 0.237
192 0.148 0.242 0.151 0.244 0.208 0.256 0.172 0.266 0.171 0.273 0.178 0.280 0.159 0.253 0.162 0.271 0.151 0.245 0.152 0.249
ECL 336 0.164 0.259 0.169 0.265 0.255 0.290 0.176 0.271 0.190 0.291 0.198 0.302 0.177 0.270 0.175 0.279 0.168 0.262 0.169 0.267
720 0.203 0.293 0.201 0.291 0.405 0.397 0.211 0.303 0.226 0.324 0.233 0.344 0.216 0.303 0.207 0.306 0.205 0.294 0.233 0.344
Avg 0.161 0.255 0.163 0.257 0.256 0.289 0.176 0.270 0.185 0.287 0.173 0.266 0.173 0.268 0.170 0.275 0.164 0.258 0.177 0.274
96 0.165 0.220 0.177 0.235 —— —— 0.197 0.244 0.202 0.222 —— —— 0.179 0.220 0.224 0.289 0.183 0.265 0.191 0.256
192 0.183 0.235 0.190 0.249 —— —— 0.230 0.274 0.216 0.240 —— —— 0.198 0.236 0.244 0.289 0.205 0.283 0.211 0.273
Solar 336 0.199 0.248 0.199 0.262 —— —— 0.244 0.281 0.236 0.255 —— —— 0.213 0.252 0.225 0.291 0.224 0.299 0.228 0.287
720 0.219 0.266 0.209 0.276 —— —— 0.247 0.280 0.252 0.264 —— —— 0.239 0.277 0.243 0.301 0.239 0.316 0.236 0.295
Avg 0.192 0.242 0.194 0.256 —— —— 0.230 0.270 0.227 0.276 —— —— 0.207 0.246 0.234 0.293 0.213 0.291 0.217 0.278
@@
Ablation on LLM-based SOTA methods. The modal disparity between time-series and language data is a challenging problem. To address this issue, we integrated the TSCC and the Time-Adapter as a plugin into the Time-LLM (Jin et al. 2024), AutoTimes (Liu et al. 2024b), and TimeCMA (Liu et al. 2025). We analyzed their impact on algorithm performance using the ETTh and ETTm datasets. Our experiments were conducted based on the parameters provided in the source code. All other experimental parameters remained unchanged when incorporating the TSCC and Time-Adapter, with the GPT-2 model consistently used as the LLM. The experimental results, listed in Table 6, indicate that errors were reduced across the board. However, Time-LLM yielded negligible performance improvements on the ETTm dataset. Notably, for the TimeCMA model, the Mean Squared Error (MSE) reduces significantly. Specifically, on the ETTh2 dataset, the MSE was reduced by 8.9%. TimeCMA uses large language models as the output of time-series text prompts. The LLM does not participate in the training, making conducting ablation experiments on Time-Adapter impossible. The experiment results highlight that our research addresses a common issue in existing studies.

### v3

@@ Introduction
Most existing LLM-based methods mainly focus on token-level alignment (Hu et al. 2025), as shown in Fig. 1 (a) (Pan et al. 2024a; Liu et al. 2025). The joint space serves as an implicit prompt, guiding the LLM in conducting time series forecasting. Semantic tokenization eliminates the dependency on auxiliary textual descriptions. However, this token-level alignment overlooks temporal and channel dependencies, making it difficult to capture dynamic temporal patterns. To address these limitations, our analysis identifies a critical oversight in existing LLM-based approaches (Jin et al. 2024; Liu et al. 2024b; Chang, Peng, and Chen 2025; Zhou et al. 2023). Regarding the semantic space composed of tokens and time steps, we propose the Temporal-Semantic Cross-Correlation (TSCC) Module to obtain enhanced semantics that represent temporal patterns, emphasize anomaly and de-anomaly patterns, thereby enhancing the interpretability of the token embeddings, as shown in Fig. 1 (b). Although we provide semantic guidance with high interpretability from LLMs, there exists an inherent difference between LLMs’ understanding of linguistic knowledge and temporal pattern. We propose to freeze the LLM and introduce an adapter to transform the model’s semantic information into the capability of modeling dynamic changes in time series, as shown in Fig. 1 (c).
In summary, we propose Semantic-Enhanced LLM, a novel framework that enhances LLM for time series forecasting with two key components: the Temporal-Semantic Cross-Correlation (TSCC) module and the Time-Adapter architecture. The TSCC captures cross-correlation of temporal embedding and semantic space with temporal dependencies and channel-wise relationships, which embed into token embedding to adapt LLM to temporal features. The Time-Adapter enhances forecasting capability by overcoming the limitation of LLMs in modeling both short-term and long-term dependencies based on historical time segments. This enables pretrained LLMs to forecast time series effectively. Our main contributions are four-fold:
@@ Methods
Figure 2: (a) An overview of SE-LLM. Red indicates trainable, while blue indicates frozen. (b) The overall framework of TSCC, with detailed processes referring to Fig. 3. (c) For the Transformer-based LLM, the Time-Adapter is embedded within the key and value vectors of the multi-head attention mechanism. JS is an abbreviation for Joint Space.
Semantic Space. The semantic space is the semantic embedding obtained by mapping the word embedding of a large language model through a linear layer, which we collectively refer to as the semantic space. Specifically, we map the Word Embedding layer from $W\in\mathbb{R}^{V\times C}$ to $S\in\mathbb{R}^{K\times C}$ (Fang et al. 2025). The pre-trained word embeddings from LLM are used as general features, and a linear transformation is applied to transfer the features to a new semantic space, allowing the model to match the time series of the TS Embeddings.
We input both the semantic space and temporal features into the TSCC to obtain enhanced semantics, as shown in Fig. 2(b). Following this, the LLM embeds Time-Adapter into the Multi-Head Attention to further refine the model’s ability to understand temporal patterns in the time series data, as shown in Fig. 2(c). This component enhances the overall temporal pattern recognition by integrating additional contextual information, enabling better forecasting performance. Finally, the Decoder component decodes the output from the LLM to produce the final prediction, which can be formulated as:
@@
where $\mathbf{Y}$ is the LLM’s output.
@@
Anomaly Pattern Modeling. In time series analysis, the anomaly noise in non-stationarity information lead to inaccurate model parameter estimation and increased prediction bias. We employ a Anomaly Modeling Variational Autoencoder (AM-VAE) (Cemgil et al. 2020) to reconstruct the mean and variance in a joint space, comprehensively capturing the underlying stochastic process of anomaly noise in temporal data, as described in Algorithm 1. The reconstruction process eliminates the reconstruction loss function, enabling the model to allow the latent variable distribution to dynamically adapt over time, thereby capturing temporal pattern.
@@
Temporal Pattern Infusion. The cross-correlation matrix $\mathbf{M}$ is computed via standardization over both temporal and semantic features, which can be formulated as:
@@
Followed by a top-K (Liu, Zeng, and Ren 2024) filtering mechanism to select the most strongly correlated positions and weighted to the output of AM-VAE to endow both anomaly and de-anomaly semantics with a temporal features that exhibits strong correlations. It can be formulated as:
@@
where $\mathbf{\overline{DC}}$ represents enhanced anomaly semantics reconstructed by the AM-VAE module from the joint space, $\mathbf{\overline{DA}}$ represents enhanced de-anomaly semantics, $\mathcal{P}$ represents the matrix in the semantic space corresponding to the indices of the top-k positions in the cross-correlation matrix.
Channel Dependency Enhancement. The aforementioned enhanced semantics are learned by integrating the temporal features TS. To prevent the loss of feature information, we further acquire channel dependencies based on the enhanced semantics modeled along the temporal dimension. Specifically, we concencate the TS Embeddings and the $\overline{\textbf{DA}},\overline{\textbf{DC}}$, that share information across channel dimensions. We employ a Multi-Layer Perceptron (MLP) for attention extraction, perform channel enhancement on the enhanced semantics in the channel dimension, which can be formulated as :
@@
$\mathbf{GA}=\mathcal{F}_{llm}(\mathrm{Attn_{a}}\odot\mathrm{H}+(1-\mathrm{Attn_{a}})\odot\mathrm{C}),$ (8)
where, $\mathcal{F}_{llm}$ is a linear layer that maps the enhanced semantics to the token space of the LLM. The resulting embeddings are enriched with semantic information that reflects temporal patterns. Finally, the de-anomalization ($\mathbf{GA}\in\mathbb{R}^{B\times N\times C}$) and anomaly ($\mathbf{GC}\in\mathbb{R}^{B\times N\times C}$) representations are mapped through a linear layer for alignment and fused to generate a unified semantic representation, which is more interpretable and suitable for LLMs analysis. It can be formulated as:
$\mathbf{Y}=\mathbf{LLM}(\mathrm{GA}+\mathrm{GC}).$ (9)
Algorithm 1 Anomaly Noise Modeling via AM-VAE Module
0: Cross-attention output: Joint Space $\mathbf{C}\in\mathbb{R}^{B\times N\times C}$
0: Decomposed anomaly semantic $\mathbf{DC}$ and de-anomalized semantic $\mathbf{DA}$ with temporal pattern
2:Latent Distribution Estimation: $\boldsymbol{\mu},\log\boldsymbol{\sigma}^{2}\leftarrow\mathcal{F}_{\mu}(\mathbf{V}),\mathcal{F}_{\sigma}(\mathbf{V}),$ where $\mathcal{F}_{\mu},\mathcal{F}_{\sigma}$: Linear mapping from $\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$.
3:Reparameterization Trick: Sample from latent space, $\mathbf{z}\leftarrow\boldsymbol{\mu}+\boldsymbol{\epsilon}\odot\exp\left(0.5\cdot\log\boldsymbol{\sigma}^{2}\right)\leftrightarrow z=\mu+\epsilon\odot\sigma,\quad\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$.
4:Anomaly Sample: In time series forecasting, the true observations are modeled as: $\mathbf{b_{t}}=f(a_{t})+\eta$, where $f(a_{t})$ is a prediction function based on historical data, and $\eta$ is a random noise term. We assume that future values follow a certain distribution, for example: $\mathbf{b_{t}}\sim(\mu_{t},\sigma^{2}_{t})$. The prediction not a single point estimate, but a probability distribution. To generate possible future trajectories, sampling from this distribution is required: $\mathbf{b_{t}}=\mu_{t}+\epsilon\cdot\sigma_{t}$, where $\epsilon~\sim N(0,1)$. The statistical characteristics of the noise may vary over time, which corresponds precisely to the per-timestep application of the Reparameterization Trick.
5:Anomaly Semantic Modeling: Decode latent variable to recover anomaly semantic, $\mathbf{DC}\leftarrow\mathcal{F}_{d}(\mathbf{z}),\quad\mathcal{F}_{d}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{C}$
6:Anomaly Decomposition: $\mathbf{DA}=\mathbf{C}-\mathbf{DC},\quad\mathbf{DC}\in\mathbb{R}^{B\times N\times C},\mathbf{DA}\in\mathbb{R}^{B\times N\times C}$, where $\mathbf{DC}$ represents the latent anomaly noise semantic in joint space, while $\mathbf{DA}$ represents the joint space where anomalous semantic are removed.
@@
Models SE-LLM TimeXier++ TimeMOE TQNet Time-CMA LLM-TS AutoTimes-GPT2 Time-LLM iTransformer
2025 2025 2025 2025 2025 2024 2024 2024
Datasets Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.352 0.393 0.375 0.404 0.349 0.379 0.379 0.403 0.366 0.396 0.403 0.420 0.360 0.397 0.398 0.410 0.387 0.418
192 0.378 0.411 0.415 0.436 0.395 0.413 0.417 0.430 0.401 0.420 0.440 0.441 0.391 0.419 0.451 0.440 0.416 0.437
336 0.391 0.420 0.442 0.451 0.447 0.453 0.452 0.457 0.412 0.431 0.471 0.457 0.408 0.432 0.473 0.451 0.434 0.450
720 0.402 0.437 0.482 0.482 0.457 0.462 0.505 0.510 0.406 0.427 0.503 0.487 0.429 0.452 0.469 0.470 0.447 0.473
avg 0.381 0.415 0.429 0.443 0.412 0.426 0.438 0.450 0.396 0.419 0.454 0.451 0.397 0.425 0.448 0.443 0.421 0.445
Weather 96 0.150 0.200 0.150 0.202 0.198 0.288 0.149 0.202 0.167 0.211 0.166 0.217 0.158 0.208 0.149 0.200 0.174 0.225
192 0.194 0.243 0.196 0.244 0.235 0.312 0.201 0.249 0.212 0.253 0.229 0.269 0.207 0.264 0.195 0.243 0.227 0.268
336 0.247 0.285 0.248 0.285 0.293 0.348 0.247 0.287 0.270 0.2927 0.278 0.302 0.262 0.298 0.245 0.282 0.290 0.309
720 0.323 0.339 0.316 0.332 0.427 0.428 0.312 0.332 0.350 0.348 0.354 0.351 0.342 0.353 0.318 0.338 0.374 0.360
avg 0.229 0.267 0.228 0.266 0.288 0.344 0.227 0.268 0.250 0.276 0.257 0.285 0.242 0.281 0.227 0.266 0.266 0.291
Traffic 96 0.350 0.243 0.367 0.266 —— —— 0.366 0.255 —— —— 0.587 0.315 0.369 0.257 0.373 0.280 0.363 0.259
192 0.373 0.254 0.379 0.364 —— —— 0.392 0.276 —— —— 0.612 0.326 0.394 0.268 0.390 0.288 0.383 0.267
336 0.390 0.263 0.417 0.301 —— —— 0.400 0.280 —— —— 0.634 0.338 0.413 0.278 0.407 0.299 0.396 0.275
720 0.429 0.285 0.453 0.314 —— —— 0.442 0.300 —— —— 0.640 0.351 0.449 0.299 0.438 0.310 0.435 0.296
avg 0.386 0.261 0.404 0.311 —— —— 0.400 0.278 —— —— 0.618 0.333 0.406 0.276 0.402 0.294 0.394 0.274
ECL 96 0.130 0.225 0.132 0.226 0.157 0.211 0.131 0.228 0.151 0.260 0.167 0.271 0.140 0.236 0.137 0.244 0.133 0.229
192 0.148 0.242 0.151 0.244 0.208 0.256 0.153 0.251 0.171 0.273 0.178 0.280 0.159 0.253 0.162 0.271 0.151 0.245
336 0.164 0.259 0.169 0.265 0.255 0.290 0.169 0.268 0.190 0.291 0.198 0.302 0.177 0.270 0.175 0.279 0.168 0.262
720 0.203 0.293 0.201 0.291 0.405 0.397 0.192 0.287 0.226 0.324 0.233 0.344 0.216 0.303 0.207 0.306 0.205 0.294
avg 0.161 0.255 0.163 0.257 0.256 0.289 0.161 0.259 0.185 0.287 0.173 0.266 0.173 0.268 0.170 0.275 0.164 0.258
Solar 96 0.165 0.220 0.177 0.235 —— —— 0.197 0.244 0.202 0.222 —— —— 0.179 0.220 0.224 0.289 0.183 0.265
192 0.183 0.235 0.190 0.249 —— —— 0.230 0.274 0.216 0.240 —— —— 0.198 0.236 0.244 0.289 0.205 0.283
336 0.199 0.248 0.199 0.262 —— —— 0.244 0.281 0.236 0.255 —— —— 0.213 0.252 0.225 0.291 0.224 0.299
720 0.219 0.266 0.209 0.276 —— —— 0.247 0.280 0.252 0.264 —— —— 0.239 0.277 0.243 0.301 0.239 0.316
avg 0.192 0.242 0.194 0.256 —— —— 0.230 0.270 0.227 0.276 —— —— 0.207 0.246 0.234 0.293 0.213 0.291
@@
Ablation on LLM-based SOTA methods. The modal disparity between time-series and language data is a challenging problem. To address this issue, we integrated the TSCC and the Time-Adapter as a plugin into the Time-LLM (Jin et al. 2024), AutoTimes (Liu et al. 2024b), and TimeCMA (Liu et al. 2025). We analyzed their impact on algorithm performance using the ETTh and ETTm datasets. Our experiments were conducted based on the parameters provided in the source code. All other experimental parameters remained unchanged when incorporating the TSCC and Time-Adapter, with the GPT-2 model consistently used as the LLM. The experimental results, listed in Table 6, indicate that errors were reduced across the board. However, Time-LLM yielded negligible performance improvements on the ETTm dataset.
Notably, for the TimeCMA model, the Mean Squared Error (MSE) reduces significantly. Specifically, on the ETTh2 dataset, the MSE was reduced by 8.9%. TimeCMA uses large language models as the output of time-series text prompts. The LLM does not participate in the training, making conducting ablation experiments on Time-Adapter impossible. The experiment results highlight that our research addresses a common issue in existing studies.